OpenStudy (newarf):

multiple (2 * 10^4) (3 * 10^3)

OpenStudy (welshfella):

first multiply the 2 numbers 2 and 3. Then calculate 10^4 * 10^3 using the laws of exponents

OpenStudy (welshfella):

0 to give 6 * 10^n what would be the value of n?

OpenStudy (newarf):

so 2000 and 3000

OpenStudy (welshfella):

????? 2 8 3 = what

OpenStudy (welshfella):

* 2 * 3

OpenStudy (newarf):

the answer would be 6*10^7 right

OpenStudy (mrnood):

You can write it as (2*3) * (10^4*10^3)

OpenStudy (welshfella):

lets perform a multiply with 2 exponential terms 10^3 * 10^5 = 10^(3+5) = 10^8 you ADD the exponents

OpenStudy (welshfella):

yes 6*10^7 is corect